OpenStudy (anonymous):

How many degrees are in (3π/2) radians?

hartnn (hartnn):

there are 180 degrees in \(\pi \) radians. so, there will be \(180/\pi \) degrees in one radians so \((3\pi/2)\times (180/\pi)=...\) just simplify this :)

OpenStudy (anonymous):

To convert angle in degree to radian it should be multiplied by π/180. If you want to convert from radian to degree the value should be multiplied by 180/π π is nothing but our good old π itself. That is approximately 3.14.

OpenStudy (anonymous):

okay than if i have a reverse question that says how many radians are in 135 degrees how do i do that @hartnn @thaahoo

OpenStudy (anonymous):

To convert angle in degree to radian it should be multiplied by π/180.

hartnn (hartnn):

simple, 180 degrees in pi radians so, there will be pi/180 radians in 1 degree so, \(135 \times \pi/180\)

OpenStudy (anonymous):

so my answer would be \[\frac{ 3\Pi }{ 4 }\]

hartnn (hartnn):

yes! correct :)

OpenStudy (anonymous):

awesome! you wouldn't happen to be good at word problems/ sine would you?

hartnn (hartnn):

there are many experts here, who are good at those. make a new post for new question, so that others can help you too :)

OpenStudy (anonymous):

lol thanks